### v1

# Align Attention Heads Before Merging Them: An Effective Way for Converting MHA to GQA

Abstract
Large language models have been shown to perform well on a variety of natural language processing problems. However, as the model size and the input sequence’s length increase, the rapid increase of KV Cache significantly slows down inference speed. Therefore GQA model, as an alternative to MHA model, has been widely introduced into LLMs. In this work, we propose a low-cost method for pruning MHA models into GQA models with any compression ratio of key-value heads. Our method is based on $\mathit{L_{0}}$ masks to gradually remove redundant parameters. In addition, we apply orthogonal transformations to attention heads without changing the model to increase similarity between attention heads before pruning training, in order to further improve performance of the model. Our method can be compatible with rotary position embedding (RoPE), which means the model after training can be fully adapted to the mainstream standard GQA framework. Experiments demonstrate that our strategy can compress up to 87.5% of key-value heads of the LLaMA2-7B model without too much performance degradation, just achieved through supervised fine-tuning.
Align Attention Heads Before Merging Them: An Effective Way for Converting MHA to GQA
Qingyun Jin1,2   Xiaohui Song2   Feng Zhou2   Zengchang Qin1 thanks: Corresponding author. 1Beihang University, Beijing, China 2OPPO AI Center, Beijing, China {jinqingyun, zcqin}@buaa.edu.cn {songxiaohui, zhoufeng1}@oppo.com
1 Introduction
Recently, Large language models (LLMs) Radford et al. (2018); Brown et al. (2020); Ouyang et al. (2022) show remarkable performance on a variety of natural language processing tasks. However, since most LLMs are based on Transformer architecture Vaswani (2017), the expansion of the sequence length during inference results in a linear increase in the memory footprint of KV Cache, significantly slowing down model inference. Therefore, reducing the size of KV Cache is a key issue for LLMs.
Multi-query attention (MQA) Shazeer (2019) and grouped-query attention (GQA) Ainslie et al. (2023) reduce KV Cache by allowing multiple attention heads to share a single key-value head, which is simple and effective. Since GQA has better inference stability and performance, it has been widely used in LLaMA 2 Touvron et al. (2023), LLaMA 3 Dubey et al. (2024), Qwen2 Yang et al. (2024), Mistral Jiang et al. (2023) and other LLMs Liu et al. (2024); Zhang et al. (2024).
A method for converting MHA model to GQA model is also proposed in Ainslie et al. (2023). They construct each group key and value head by mean-pooling all the original heads within that group and then uptrain the model to restore the model performance. However, the computational resources required for uptraining are unaffordable in most conditions.
Figure 1: An illustration of our pruning training process. The training loss consists of two parts: distillation loss between teacher model and pruning model, and pruning loss to impose key-value projection matrices gradually transferred to new ones. After pruning, original key-value projection matrices will be discarded, then we get a standard GQA model.
In this work, we use $\mathit{L_{0}}$ masks Louizos et al. (2017) to transfer original key-value projection matrices to new ones. Figure 1 shows our concept. In addition, based on the idea of computational invariance in the model Ashkboos et al. (2024), we apply orthogonal transformations to the matrices of attention heads without changing the model before pruning training: We measure the similarity between KV Caches of different attention heads after orthogonal transformation, then divide attention heads with high similarity of KV Caches into the same group. The similarity among the KV Caches derived from the attention heads within the same group is then maximized through an orthogonal transformation. Experiments show that this method can significantly improve the performance of the pruned model after the transformation of attention heads. Yu et al. (2024) is the most relevant work to ours; however, their method is not fully compatible with RoPE Su et al. (2024). Our contributions are as follows.
  • We propose a general and low-cost method for converting MHA structure to GQA, which can compress the key and value heads to any percentage and basically restore performance after supervised fine-tuning.
  • Our study provides a new perspective for evaluating the similarity between attention heads, which presents new insights for future research related to compressing KV Cache.
  • We conduct experiments on pruning training the LLaMA2-7B model Touvron et al. (2023) into GQA-16, GQA-8, and GQA-4, separately. The model performance does not decrease significantly compared to that of the full-size model.
    Figure 2: In each block of LLaMA2-7B, the average cosine similarity is calculated between every two key and value caches. For convenience, the average similarities shown in this figure are their absolute values. It can be seen that most pairs of KV Caches are almost orthogonal. This may explain why directly merging key-value heads causes significant loss.
2 Related Works
2.1 $\mathit{L_{0}}$ regularization
$\mathit{L_{0}}$ regularization Louizos et al. (2017) is a structured pruning approach that transforms a pruning problem into an optimization problem under constraints. The pruning process is performed simultaneously with model optimization by introducing trainable masks. With the wide application of LLMs, this method has been applied to compressing LLMs. In the work of Wang et al. (2019), the $\mathit{L_{0}}$ method is applied based on low-rank pruning to achieve further improvements in effectiveness, and they propose to gradually increase the target size at a linear rate during the process of pruning training. In CoFi Xia et al. (2022), the $\mathit{L_{0}}$ method is applied directly to LLMs by introducing pruning masks with different granularities. They prune the hidden dimension, the intermediate dimension, the number of attention heads, and even an entire MHA or FFN layer. The subsequent work Sheared-LLaMA Xia et al. (2023) incorporates previous methods and specifies the target structure so that the pruned model can be directly adapted to standard LLM frameworks.
2.2 Transfer MHA to GQA
Ainslie et al. (2023) proposes GQA for the first time, in which MHA is converted to GQA using mean pooling initialization. However, this method requires uptraining to restore performance and incurs significant computational costs. Yu et al. (2024) keeps the corresponding parameters based on the principal components of the collected KV Caches, then uses LoRA Hu et al. (2021) to fine-tune the model to restore performance. Chen et al. (2024) proposes to regroup attention heads based on the criterion of cosine similarity and allows for varying group sizes. However, none of the aforementioned improvement methods can be fully adapted to the standard GQA model.
2.3 Compressing model based on the principal components of features
Some previous works Liu et al. (2023); Yu and Wu (2023) have pointed out that the features of LLMs are generally low-rank. Therefore, identifying and deleting the low-rank components of the model is an effective method for model compression.
Low Rank BERT Noach and Goldberg (2020) reduces the number of parameters and increases inference speed by decomposing the weight matrices into two low-rank matrices. SliceGPT Ashkboos et al. (2024) introduces the idea of computational invariance in Transformer architecture and removes columns or rows of the transformed weight matrices to reduce model size. Yu et al. (2024) applies orthogonal transformations to key-value projection matrices by analyzing the low-rank features of KV Cache.
3 Method
In this section, we will specifically describe our method. Our method consists of two parts, transformation of the network and pruning training. Transformation of the network represents applying orthogonal transformations to the projection matrices in order to increase the similarity between attention heads of the same group, so that we can increase efficiency of model optimization. The pruning training process combines pruning using $\mathit{L_{0}}$ masks Louizos et al. (2017) and knowledge distillation Gou et al. (2021).
3.1 Motivation
To analyze the characteristics of KV Cache, we follow a prior calibration method for LLMs Frantar and Alistarh (2023); Sun et al. (2023) in order to obtain calibration data: Sample 128 sequences from the C4 Raffel et al. (2020) training set and each sequence is 2048 tokens long, 262144 tokens in total. Then perform model inference on LLaMA2-7B and collect KV Caches, i.e.,
$K=[K_{1};\ldots;K_{H}]\quad\quad V=[V_{1};\ldots;V_{H}]$ (1)
where $K,V\in\mathbb{R}^{d\times N}$ are KV caches corresponding to each block, which can be divided into $K_{i},V_{i}\in\mathbb{R}^{d_{H}\times N}$, $N$ is the number of tokens, ${d}$ is embedding dimension and ${H}$ represents the number of heads in each MHA, ${d_{H}}$ is set to ${d/H}$, then we can calculate the average cosine similarity between each of two heads as follows:
$SimK_{i,j}^{ori}=\frac{1}{N}\sum_{n=0}^{N-1}cos({{K}_{i}[n]\cdot{K}_{j}[n]})$ (2)
$SimV_{i,j}^{ori}=\frac{1}{N}\sum_{n=0}^{N-1}cos({{V}_{i}[n]\cdot{V}_{j}[n]})$ (3)
where $i,j$ are any two of attention heads in the same block, $n$ represents the $n^{th}$ token in this cache. As shown in Figure 2, we notice that while a few pairs of KV Caches share high cosine similarity with each other, the vast majority of them are almost orthogonal. This is the reason why directly mean-pooling projection matrices results in significant loss so that uptraining is needed to restore performance.
However, according to Yu et al. (2024), KV Caches are Low-rank. Given that these caches occupy only a portion of spatial dimensions, applying appropriate orthogonal transformations to the projection matrices to align key and value caches can reduce the difficulty of model optimization. Fortunately, this approach is feasible.
3.2 Preliminaries
Given two sets of vectors of the same shape: $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}\in\mathbb{R}^{M\times N}$ and $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{N}\}\in\mathbb{R}^{M\times N}$, how to find the optimal orthogonal transformation that aligns the two sets of vectors? This kind of problems is called Orthogonal Procrustes problem, and its mathematical formulation is as follows:
$\min_{\mathbf{Q}}\|\mathbf{Q}\mathbf{X}-\mathbf{Y}\|_{F}^{2}$ (4)
The optimal orthogonal transformation can be derived from SVD of the matrix $\mathbf{Y}\mathbf{X}^{T}$, the general solution is Schönemann (1966):
Perform SVD on the covariance matrix of X and Y,
$(\mathbf{Y}\mathbf{X}^{T})=\mathbf{\Phi}\mathbf{\Sigma}\mathbf{\Psi}^{T}$ (5)
Then obtain the optimal orthogonal matrix $\mathbf{Q}$:
$\mathbf{Q}=\mathbf{\Psi}\mathbf{\Phi}^{T}$ (6)
We can use the same way to align any two KV Caches from different attention heads in the same block. Furthermore, if we want to align more than two sets of caches, Generalized Procrustes analysis Wikipedia contributors (2022) is a good solution. The detailed description is shown in algorithm 1.
Algorithm 1 Generalized Procrustes Analysis
0:  Matrices $X_{1},X_{2},\ldots,X_{H}$
0:  Aligned matrices $Y_{1},Y_{2},\ldots,Y_{H}$
  Initialize $Y_{i}=X_{i}$ for all $i$
  Compute mean shape $\bar{M}=\frac{1}{H}\sum_{i=1}^{H}Y_{i}$
  repeat
     for $i=1$ to $H$ do
        Compute $\Phi_{i}\Sigma_{i}\Psi_{i}^{T}=\text{SVD}(Y_{i}^{T}M)$
        Update $Y_{i}=Y_{i}\Psi_{i}\Phi_{i}^{T}$
     end for
     Update mean shape $\bar{M}=\frac{1}{H}\sum_{i=1}^{H}Y_{i}$
  until convergence
  return  $Y_{1},Y_{2},\ldots,Y_{H}$
Figure 3: This figure shows the average cosine similarity of key and value caches between any two heads before and after an applying transformation in some blocks of LLaMA2-7B. Appropriate orthogonal transformations can significantly improve the cosine similarity between KV Caches.
3.3 Transformation of attention heads
To calculate the optimal orthogonal matrix for each pair of key and value heads, we collect KV Caches according to the method mentioned above. Here, we use two criteria to perform the calculations.
Based on cosine similarity. Firstly normalize each vector in $K_{i}$ and $V_{i}$ to roughly reduce the influence of magnitude of the vector:
$\hat{K}_{i}[*]=\frac{K_{i}[*]}{\|K_{i}[*]\|}$ (7)
$\hat{V}_{i}[*]=\frac{V_{i}[*]}{\|V_{i}[*]\|}$ (8)
then we can get the optimal orthogonal matrix $Q_{V}$ to align any two value caches, taking $\hat{V_{i}}$ and $\hat{V_{j}}$ as example:
$(\hat{V_{i}}\hat{V_{j}}^{T})=\Phi\Sigma\Psi^{T}$ (9)
$Q_{V_{j}}=\Psi\Phi^{T}$ (10)
For each block, the output of self-attention layer can be seen as the sum of all attention heads:
$\text{MultiHead}(W_{Q},W_{K},W_{V},W_{O})$ (11)
$=\sum_{i=1}^{H}(W_{O_{i}}(W_{V_{i}}X)\text{Softmax}\left(\frac{(W_{K_{i}}X)^{T}(W_{Q_{i}}x)}{\sqrt{d_{H}}}\right))$
where the projection matrices in the attention heads are $W_{Q_{i}},W_{K_{i}},W_{V_{i}}\in\mathbb{R}^{d_{H}\times d}$ and $W_{O_{i}}\in\mathbb{R}^{d\times d_{H}}$, $X\in\mathbb{R}^{d\times len}$ represents the input sequence, and $x\in\mathbb{R}^{d\times 1}$ represents the current token. For brevity, RoPE is ignored here. Then we can fuse the orthogonal matrix into the value projection matrix ${W}_{V_{j}}$ and the output projection matrix $W_{O_{j}}$ to ensure computational invariance:
$W^{{}^{\prime}}_{V_{j}}=Q_{V_{j}}W_{V_{j}}$ (12)
$W^{{}^{\prime}}_{O_{j}}=W_{O_{j}}Q_{V_{j}}^{T}$ (13)
As for $W_{Q}$ and $W_{K}$, due to the existence of RoPE, orthogonal transformation cannot be applied directly. However, we can divide the d-dimension space into $d/2$ sub-spaces and apply the orthogonal transformation in every two dimension just like RoPE, which is to say the orthogonal matrix for key projection matrix should be in this form:
$R_{K_{j}}=\begin{pmatrix}R_{\theta_{1}}&0&\cdots&0\\ 0&R_{\theta_{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&R_{\theta_{d/2}}\end{pmatrix}$ (14)
where $R_{\theta_{\cdot}}$ is a 2D rotation matrix. Then we fuse the orthogonal matrix $R_{K_{j}}$ into the query projection matrix $W_{Q_{j}}$ and key projection matrix $W_{K_{j}}$:
$W^{{}^{\prime}}_{Q_{j}}=R_{k_{j}}W_{Q_{j}}$ (15)
$W^{{}^{\prime}}_{K_{j}}=R_{k_{j}}W_{K_{j}}$ (16)
So, we get
$q_{s}^{{}^{\prime}T}k^{{}^{\prime}}_{t}$ $=(R_{\Theta,s}^{d}(R_{K_{j}}{W}_{Q_{j}}){x}_{s})^{T}({R}_{\Theta,t}^{d}(R_{K_{j}}{W}_{K_{j}}){x}_{t})$ (17)
$={x}_{s}^{T}{W}_{Q_{j}}^{T}(R_{K_{j}}^{T}{R}_{\Theta,t-s}^{d}R_{K_{j}}){W}_{K_{j}}{x}_{t}$
$={x}_{s}^{T}{W}_{Q_{j}}^{T}{R}_{\Theta,t-s}^{d}{W}_{K_{j}}{x}_{t}$
$=(R_{\Theta,s}^{d}{W}_{Q_{j}}{x}_{s})^{T}({R}_{\Theta,t}^{d}{W}_{K_{j}}{x}_{t})$
$={q}_{s}^{T}{k}_{t}$
where $q_{s}$ represents the query of the $s^{th}$ position and $k_{t}$ represents the key of the $t^{th}$ position. This transformation doesn’t change the model either.
In this way, given any two key or value caches, we can use this method to calculate the maximum cosine similarity achievable.
$SimK_{i,j}^{after}=\frac{1}{N}\sum_{n=0}^{N-1}cos({{K}_{i}[n]\cdot(R_{K_{j}}{K}_{j}[n])})$ (18)
$SimV_{i,j}^{after}=\frac{1}{N}\sum_{n=0}^{N-1}cos({{V}_{i}[n]\cdot(Q_{V_{j}}{V}_{j}[n])})$ (19)
Noticing $SimV_{i,j}^{after}$ is equal to $SimV_{j,i}^{after}$, so is $SimK^{after}$. Figure 3 shows the cosine similarity between KV Caches before and after applying the transformation.
Based on Euclidean distance. Similar to applying transformations based on cosine similarity, we also apply transformations based on Euclidean distance between two KV Caches. In this case, we don’t normalize vectors and the similarity between two caches can be described as the negative value of the Euclidean distance of them, for brevity, only some key formulas are displayed here:
$({V_{i}}{V_{j}}^{T})=\Theta\Lambda\Omega^{T}$ (20)
$P_{V_{j}}=\Omega\Theta^{T}$ (21)
$SimV_{i,j}^{after}=-\frac{1}{N}\sum_{n=0}^{N-1}\|{V}_{i}[n]-(P_{V_{j}}{V}_{j}[n])\|_{F}$ (22)
In the next section, we will compare the performances of the two criteria.
3.4 Find better grouping method
After obtaining the similarity scores between every pair of attention heads, we can regroup attention heads based on these scores. We define the similarity score of a group as the sum of similarity scores between every pair of attention heads within that group, and the total similarity score for each grouping result is the sum of similarity scores of all groups. Our objective is to determine the grouping result with the highest total score [^1] While the highest similarity between pairs within the same group does not necessarily equate to the lowest cost in terms of converging to the same parameters during pruning, this strategy remains acceptable considering computation and time costs.. We use $SimK^{after}$ and $SimV^{after}$ as grouping criteria, respectively. In the next section, we will compare the performances of these two ways. The mathematical expression of the score of a grouping result $A=\{A_{1},A_{2},\cdots,A_{G}\}$ is as follows:
$Score_{key}(A)=\sum_{g=1}^{G}\sum_{0\leq i<j<D}SimK_{A_{g}[i],A_{g}[j]}^{after}$ (23)
$Score_{value}(A)=\sum_{g=1}^{G}\sum_{0\leq i<j<D}SimV_{A_{g}[i],A_{g}[j]}^{after}$ (24)
where $A_{g}$ is the $g^{th}$ group in $G$ groups, the elements in $A_{g}$ are the serial number of an attention head and there are $D=H/G$ heads in a group.
We use Simulated Annealing Algorithm to get the best grouping result: Exchange two random heads in different groups and calculate the new score, accepting the new result if it reaches a higher score. Repeat this process for multiple iterations. Because initialization has a significant impact on the final result, we run the algorithm multiple times. The details of the algorithm 2 are shown below.
Algorithm 2 Simulated Annealing
0:  $maxIter$, $epoch$, $SimK$ or $SimV$
0:  grouping result with the highest score $bestG_{n}$
  Set $score_{best}$ $\leftarrow$-$\infty$ - ∞
  for $i=1$ to $epoch$ do
     Initialize solution $G_{n}$ randomly
     $score_{current}$leftarrow calculate_score($G_{n}$, $SimV$)
     if $score_{current}$ $>$ $score_{best}$ then
        Set $score_{best}$leftarrow $score_{current}$
        Set $bestG_{n}$leftarrow $G_{n}$
     end if
     for $j=1$ to $maxIter$ do
        $G^{{}^{\prime}}_{n}$leftarrow exchange random elements from different groups in $G_{n}$
        $score_{new}$leftarrow calculate_score($G^{{}^{\prime}}_{n}$, $SimV$)
        if $score_{new}$ $>$ $score_{current}$ then
           Set $G_{n}$leftarrow $G^{{}^{\prime}}_{n}$
           $score_{current}$leftarrow $score_{new}$
           if $score_{new}$ $>$ $score_{best}$ then
              Set $score_{best}$leftarrow $score_{new}$
              Set $bestG_{n}$leftarrow $G_{n}$
           end if
        end if
     end for
  end for
  return  $bestG_{n}$
After grouping, we can use Generalized Procrustes analysis to align attention heads in the same group.
3.5 Adaptation of $\mathit{L_{0}}$ regularization
During pruning training, we add new projection matrices which are initialized by mean-pooling all the original heads within that group to the model Ainslie et al. (2023), here we use $\tilde{W}_{K_{k,g}}$ or $\tilde{W}_{V_{k,g}}$ to represent new projection matrices of the $g^{th}$ group in the $k^{th}$ block of the model:
$\tilde{W}_{K_{k,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{K_{k,(g-1)*D+i}}$ (25)
$\tilde{W}_{V_{k,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{V_{k,(g-1)*D+i}}$ (26)
These new matrices will be trained together with the model and replace original key-value heads after pruning. Assume the model has $N_{blocks}$ blocks and $H$ heads in an attention layer, we introduce $\mathit{L_{0}}$ masks $z\in\mathbb{R}^{N_{blocks}\times H}$ Louizos et al. (2017) to achieve this goal:
$W_{K_{k,j}}^{apply}=z_{k,j}W_{K_{k,j}}+(1-z_{k,j})\tilde{W}_{K_{k,g}}$ (27)
$W_{V_{k,j}}^{apply}=z_{k,j}W_{V_{k,j}}+(1-z_{k,j})\tilde{W}_{V_{k,g}}$ (28)
where $g=\lceil\frac{j}{D}\rceil$, $W_{K_{k,j}}$ and $W_{V_{k,j}}$ are the original projection matrices, $\tilde{W}_{K_{k,g}}$ and $\tilde{W}_{V_{k,g}}$ are the newly added projection matrices, $W_{K_{k,j}}^{apply}$ and $W_{V_{k,j}}^{apply}$ are the projection matrices employed during pruning. Following the $\mathit{L_{0}}$ regularization approach, we parametrize the pruning masks to hard concrete distributions. Initially, each mask is set $z=1$, we constrain the masks to zero during pruning Xia et al. (2023). And the original projection matrix will be transferred to the new matrix when $z=0$. Unlike traditional $\mathit{L_{0}}$ regularization, we aim to eliminate any original key or value heads and just utilize $\mathit{L_{0}}$ masks to gradually transfer the original heads to newly added heads. All masks across blocks are constrained by a single loss function:
~$\tilde{\mathcal{L}}_{L_{0}}=\left((\frac{1}{N_{block}H}\sum z)-T\right)+\left((\frac{1}{N_{block}H}\sum z)-T\right)^{2}$ (29)
where $T$ is the target size and equals zero after sparsity warm-up steps.
We use vanilla KL loss and BiLD loss Li et al. (2024) to encourage alignment of student logits with teacher logits.
~$\tilde{\mathcal{L}}_{distill}=\tilde{\mathcal{L}}_{KL}+\tilde{\mathcal{L}}_{BiLD}$ (30)
To sum up, the overall training loss is:
=~$\mathcal{L}=\tilde{\mathcal{L}}_{distill}+\tilde{\mathcal{L}}_{L_{0}}$ (31)
4 Experiments
4.1 Settings
Model configurations. We apply our method to the LLaMA2-7B model Touvron et al. (2023) throughout all experiments. We will convert the source model to GQA-16, GQA-8, and GQA-4, and compare them to the fine-tuned full-size model separately.
Datasets. We use the following open-source datasets for pruning training and evaluation: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challengeClark et al. (2018), SIQA Sap et al. (2019) and OpenbookQA Mihaylov et al. (2018). The size and instruction template of each dataset are listed in Appendix B.
Implementation Details. We use 1 NVIDIA A100 GPU to perform model transformation, and 8 NVIDIA A100 GPUs for supervised fine-tuning (SFT) the teacher model and pruning training. We randomly select 128 sequences of 2048 tokens long from the C4 training set Raffel et al. (2020) as calibration data in model transformation. In all experiments, the initial learning rate is 1e-5 for the model parameters and 1e-2 for the pruning masks. The cosine scheduler is employed to reduce the learning rate to 0 by the end of training. We perform 2 epochs of SFT on the teacher model, 5 epochs of pruning training on GQA-16, 15 epochs on GQA-8, and 20 epochs on GQA-4.
More hyperparameter settings can be found in Appendix A.
4.2 Ablation studies
We test the impact of different similarity evaluation criteria (see Section 3.3) and grouping strategies (see Section 3.4). All results are presented in Table 1. Here, "baseline" refers to pruning directly without any transformation, "default grouping" refers to merging adjacent attention heads, "grouping by key" and "grouping by value" indicate grouping attention heads based on key or value cache similarity. "cos" and "dist" represent the transformation based on cosine similarity or Euclidean distance.
4.3 Main Results
We report the experimental results in Table 1. Except for one set of experiments, all transformed models outperform the baseline. As the sparsity of key-value head increases, the advantage of model transformation becomes more obvious, demonstrating the effectiveness of aligning the attention heads. While the results of the GQA-8 don’t meet expectations, in other experiments, models transformed based on value cache similarity, using Euclidean distance as the criterion, achieved the best performance.
Model Methods BoolQ PIQA HellaSwag WinoGrande ARC-C ARC-E OpenbookQA SIQA Avg.
MHA Teacher 89.42 77.15 87.62 85.16 73.91 85.09 82.00 78.10 85.48
GQA-16 baseline 86.85 81.39 87.08 82.95 66.22 80.88 79.00 77.53 84.60
default grouping cos 88.62 81.50 89.08 83.82 67.56 81.40 81.40 77.53 86.08
dist 87.98 81.61 89.42 83.14 68.23 82.63 79.20 78.04 86.16
grouping by key cos 88.47 80.79 88.26 82.64 69.23 82.28 79.20 77.94 85.53
dist 88.44 80.25 88.96 82.72 70.23 80.35 77.40 77.28 85.68
grouping by value cos 87.61 81.23 88.39 82.79 69.23 80.35 75.80 77.12 85.28
dist 87.80 82.10 89.66 83.50 71.24 81.93 80.80 77.74 86.35
GQA-8 baseline 84.56 79.71 83.53 80.90 60.54 75.26 75.00 76.25 81.65
default grouping cos 86.76 80.52 86.66 81.06 64.88 79.30 77.60 76.92 84.01
dist 86.76 81.61 87.25 82.64 65.22 81.05 77.80 76.61 84.54
grouping by key cos 86.91 80.68 87.01 82.56 64.21 80.00 76.60 76.20 84.24
dist 86.79 80.03 86.39 82.56 65.89 80.00 78.40 76.46 83.94
grouping by value cos 86.39 76.28 81.80 79.64 63.54 74.74 69.20 73.90 80.32
dist 86.60 81.50 86.96 83.74 67.22 79.47 79.00 76.31 84.42
GQA-4 baseline 81.86 76.93 76.97 78.30 55.52 73.86 69.80 74.56 77.03
default grouping cos 85.47 78.89 83.18 81.53 59.53 77.02 74.40 75.49 81.53
dist 84.83 79.27 83.72 80.74 59.53 77.54 76.80 75.54 81.77
grouping by key cos 85.41 79.38 83.37 80.90 61.20 77.19 74.40 75.49 81.66
dist 85.26 78.73 81.14 81.21 62.54 75.79 73.20 75.18 80.37
grouping by value cos 86.18 79.38 84.05 82.16 60.87 76.67 74.00 75.44 82.17
dist 85.69 79.54 84.32 82.00 63.21 77.89 75.40 75.64 82.36
Table 1: Performances of prompt tuning on LLaMA2 models with various Methods. The last column, Avg. (Average Accuracy), indicates the average accuracy of all these sub datasets.
4.4 Analysis of the results
The experimental results indicate that the grouping of attention heads does have an impact on the performance of pruned models, and that grouping attention heads based on value cache similarities is beneficial for the model performance. This is no surprise, as the value cache directly contributes to the output vector of the attention layer. Although our experiments utilize $\mathit{L_{0}}$ masks to accelerate training process, model transformation can benefit any MHA to GQA conversion process.
Figure 4: The shared sparsity of $\mathit{L_{0}}$ masks across blocks allows different pruning speeds for different blocks, leading to a more stable training process.
In addition, during the experiment, we found that the model with introduction of new KV heads performs much better than the model retaining original ones. That’s why we choose not to retain any original KV heads, this setting also allows different pruning speeds for different blocks. Figure 4 shows the actual average size of masks in each block at different target sizes.
5 Conclusion
In this paper, we propose a general method for pruning an MHA model into a GQA model with any compression ratio of key-value heads. We find that applying appropriate orthogonal transformations to the model can increase the similarity between key-value heads without changing the model, thereby reducing the difficulty of model pruning. Furthermore, we introduce $\mathit{L_{0}}$ masks during pruning training, which reduce the impact of directly eliminating parameters on the model. Our method is applicable to all key-value head pruning conditions.
Limitations
Our work has two main limitations. First, we don’t delve into the grouping method, and current grouping method may not be the optimal one. How to find a more reasonable grouping method is one of the future research directions. Moreover, our method entirely relies on the statistical mathematical features of each attention head, without considering semantic information of each attention head. In fact, compressing attention heads based on semantic information is also a promising direction Tang et al. (2024).
Appendix A Hyperparameter settings
To reduce memory usage, we employ DeepSpeed during both SFT and pruning training, we set $k$=16 for BiLD loss Li et al. (2024). During the pruning training process, the sparsity warm-up steps account for 30% of the total steps, during which the target size of the $\mathit{L_{0}}$ masks decreases linearly to zero. The maximum pruning steps comprise 80% of the total steps, after which the mask training ceases, only the model parameters are adjusted. Some more hyperparameter settings for SFT teacher model and pruning training are shown in Table 2.
SFT teacher pruning training
batch size 128 64
micro batch size 4 1
warmup steps 16 32
initial lr of masks \ 1e-2
initial lr of model 1e-5
Table 2: Some hyperparameters setting for experiments.
Appendix B Details of datasets
The sizes of sub datasets are shown in Table  3.
datasets train test
BoolQ 9427 3270
PIQA 16113 1838
HellaSwag 39905 10042
WinoGrande 40398 1267
ARC-C 1119 299
ARC-E 2251 570
OpenbookQA 4957 500
SIQA 33410 1954
total 147580 19740
Table 3: Sizes of different datasets
The template of each dataset can be seen in Table  LABEL:tab:templates.
Table 4: The template of each dataset
Dataset Template
Arc-C Arc-E OpenbookQA Which color shirt will reflect the most light on a hot, sunny day? Choices: [’black’, ’blue’, ’red’, ’white’] Answer:
HellaSwag Please choose the most appropriate text to complete the passage below:
Passage: A male athlete puts powder on his hands. he
Choices: [’bends and inspects his hands for damage.’, ’shakes them shakingly before putting them in his mouth.’, ’mounts a high beam in the gym.’, ’then jumps up and does a high jump.’]
Answer:
BoolQ The Coroner – The BBC announced on 2 March 2017 that there would be no further series.
Question: will there be a second series of the coroner?
Choices: [’true’, ’false’]
Answer:
Winogrande Choose the most sensible text to replace the ’_’ in the following sentence: Natalie was less religous than Patricia, therefore _ attended church services more often on Sundays.
Choices: [’Natalie’, ’Patricia’]
Answer:
PIQA Goal: how do you flood a room?
Choose the most sensible solution to achieve the goal. Choices: [’fill it with objects.’, ’fill it with water.’]
Answer:
SIQA Sasha took him to vegas for a vacation.
Question: How would Sasha feel afterwards??
Choices: [’sad’, ’depressed’, ’fulfilled’]
Answer: